OpenStudy (anonymous):

~His work shown below.. to solve this probem~ Victor calculates the percentage interest rate he needs to make one hundred dollars in simple annual interest over 2 years on a deposit of one thousand dollars. Where did he go wrong? He applied the principal and interest to the wrong variables. He should have divided 2 by 12 to get the value for t. He misrepresented the earned interest with the final balance. He should have multiplied r by 100 to get the interest rate as a percentage.

OpenStudy (anonymous):

\[i = prt\] \[100 = 1,000 \times r \times2\] \[r = 0.05%\]

OpenStudy (anonymous):

@amistre64 Please help me!!

OpenStudy (anonymous):

@kelliegirl33 help!

OpenStudy (anonymous):

I = PRT 100 = (1000)(R)(2) 100 = 2000R 100/2000 = R .05 = R (.05)(100) = 5% answer : D

OpenStudy (anonymous):

@kelliegirl33 wow you understood that it was 0.005% i wasnt able to put the %. Thanks so much!

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

by the way .05 = 5% not .005%